OpenStudy (anonymous):

2. Determine which ordered pair in the form (a, b) is a solution of the given equation. 4a – 5b = 24 (Points : 1) (6, 0) (–5, –2) (0, 6) (–2, –5)

OpenStudy (anonymous):

can some one tell me how to solve it, but not solve it for me please

OpenStudy (ivettef365):

just substitute the numbers to see which one is true example lets get (0,6) (a,b) 4a - 5b = 24 4(0) - 5(6) = 24 0 - 30 = 24 - 30 = 24 not true , keep trying

OpenStudy (anonymous):

ok thank you